# The Battle of the Water Futures

### v1

@@ 2.3 System Interventions
This chapter provides a summary of all possible interventions available to participants in the BWF competition. All interventions have been introduced in previous sections. Here, we consolidate this information to give participants a complete overview of the decision space they can navigate when developing their strategies.
2.3.1 Non-Revenue Water Interventions
Participants can decide how much (as a percentage) of each water utility’s yearly budget should be spent on improving pipe infrastructure to reduce non-revenue water. Improving the pipe infrastructure means decreasing its average age, which is linked to the different leak classes and implies the amount of non-revenue water. There exist two policies that distribute a utility’s budget across its municipalities:
@@ 2.3.1 Non-Revenue Water Interventions
  • by_leak_class Uses the budget to improve the leak class (by one) of each municipality in a greedy way (worst cases first) until no budget is left.
  • by_population Distributes the budget proportionally to the municipalities’ population size.
@@ 2.4 System Requirements
2.4.1 Total Annualized Cost
  • Formula: $TAC=\sum_{j}\frac{K_{j}}{L_{j}}+OPEX+(coupon×P_{issue})$
  • Calculated over all dimensions (utility, time). Ranked based on cumulative amount across time and utilities.
  • unit: €
2.4.2 Total Annualized GHG Emissions
  • Formula: $\text{GHG}=\sum_{j}\frac{K_{j}\times EF^{\text{emb}}_{j}}{L_{j}}+\sum_{t}\frac{E_{t}\times EF}{1000}$
  • Calculated over all dimensions (utility, time). Ranked based on cumulative amount across time and utilities.
  • unit: tCO2eq
2.4.3 Service Reliability
  • Formula: $R^{\text{service}}=1-\dfrac{\sum_{i,t}U_{i,t}}{\sum_{i,t}D_{i,t}}$
  • Calculated over all dimensions (municipality, household class, and time). Participants will be ranked based on one specific combination of these dimensions, which is kept uncertain and can change between stages.
  • unit: dimensionless between 0 and 1 (higher is better)
2.4.4 Affordability Fairness
  • Formula: $AF=\frac{p_{v}×V_{lifeline}+F_{fixed}}{I_{p20}}×100$
  • Calculated over all dimensions (municipality, household class, and time). Participants will be ranked based on one specific combination of these dimensions, which is kept uncertain and can change between stages.
  • unit: % (lower is better)
This section defines all variables and notation used in the KPI formulas.
Symbol Units Description
General / Indices
i Index of node or municipality
t Time index (e.g., hourly step)
j Index of intervention
N - Total number of nodes
T hours Total number of simulation hours in the evaluation period
Demand Variables
$D_{i,t}$ m³/h Water demand at node i and time t
Qi,tQ{i,t} m³/h Delivered volume at node i and time t
$U_{i,t}=\max(D_{i,t}-Q_{i,t},0)$ m³/h Undelivered demand at node $i$, time $t$
Financial Variables
$K_{j}$ Capital expenditure (CAPEX) for intervention j
$L_{j}$ yr Asset lifetime of intervention j
OPEX €/yr Operating expenditures (energy, maintenance, etc)
$r_{f}$ Risk-free rate (e.g., 0.03 = 3 %)
cs Base credit spread (e.g., 0.01 = 1 %)
a Spread sensitivity to demand (e.g., 0.02 = 2 %)
d Random investor demand factor $\in$ [0.8, 1.2]
coupon Coupon rate of issued bond = $r_{f}+cs+a(1–d)$
$P_{\text{issue}}$ Principal amount of issued bond
Environmental Variables
$E_{t}$ kWh Pumping energy at time t
EFEF kg CO2/kWh Electricity emission factor
$EF^{emb}_{j}$ t CO2 / € or unit Embedded emission factor for intervention j
Socio-Economic Variables
$p_{v}$ €/m³ Volumetric tariff
$F_{\text{fixed}}$ €/month Fixed charge per household
$V_{\text{lifeline}}$ m³/month Lifeline volume (e.g., 1.5 m³/person/month)
$I_{p20}$ €/month 20th-percentile monthly income across nodes
@@ Acknowledgements
Thanks to everyone.
@@
Reference to the Zenodo repo.
Mention the data providers and the licenses.

### v4

@@ 2.3 System Levers
This chapter provides a summary of all the options available to participants in the BWF competition (they have all been introduced in previous sections already).
Here, we consolidate this information to give participants a complete overview of the decision space they can navigate when developing their masterplan.
2.3.1 Policies
Policies encompass regulatory and operational rules, such as pricing structures, budget allocations, and maintenance protocols. Once set, policies remain in effect until explicitly amended.
2.3.1.1 National Budget Allocation (National)
Participants must decide the strategy to allocate the national budget across the water utilities. The policy can be a predefined one or follow a custom allocation of the funds:
  • by_population: Allocate the funds according to each province population.
  • by_inverse_population: Allocate the funds to provinces with less population (less revenue).
  • by_income: Allocate the funds according to each province economic activity.
  • by_inverse_income: Allocate the funds to provinces with less economic activity (less revenue).
  • custom: Allocate the budget according to the specified share for each water utility (must sum to 1).
year: 2025 national_policies: budget_allocation: policy: by_population # or "by_income" or the corresponding inverse policies
year: 2026 national_policies: budget_allocation: policy: custom policy_args: WU01: 0.12 WU02: 0.25 # and so on...
2.3.1.2 Non-Revenue Water Mitigation Budget (Utility)
Participants must decide each water utility’s yearly budget used to mitigate non-revenue water (NRW). This budget is used to improve the municipalities innner distribution network (IDN). More precisely, this budget descreases the municipalities IDN’s average age, which in turn improves the NRW class of the municipality, leading to a reduction of the NRW component. The policy can be a predefined one or follow a custom allocation of the funds:
@@ 2.3.1.2 Non-Revenue Water Mitigation Budget (Utility)
  • by_nrw_class: Allocate the budget to improve by one NRW-class each municipality in a greedy way (worst cases first) until no budget is left in that year.
  • by_population: Allocate the budget according to each municipality population.
  • custom: Allocate the budget according to the specified share for each municipality (must sum to 1).
year: 2025 water_utility: WU01 policies: nrw_mitigation: budget: 30000 policy: by_nrw_class # or "by_population"
year: 2026 water_utility: WU01 policies: nrw_mitigation: budget: 30000 policy: custom policy_args: GM0001: 0.02 GM0002: 0.02 # and so on...
2.3.1.3 Water Pricing (Utility)
Participants must decide the water pricing adjustment strategy for each year. They have two options: increase all water price components according to inflation, or define a custom policy by specifying the percentage increase for each component (e.g., 2%).
year: 2025 water_utility: WU01 policies: pricing_adjustment: policy: by_inflation
year: 2026 water_utility: WU01 policies: pricing_adjustment: policy: custom # Specify the percentage increase for each pricing component policy_args: fixed_component: 0.03 # Annual increase for fixed costs (3%) variable_component: 0.02 # Annual increase for variable costs (2%) selling_price: 0.05 # Annual increase for water sales to other provinces (5%)
2.3.1.4 Bond Issuance (Utility)
Whenever the water utility is unable to cover its expenditures in a specific year, it finances the resulting deficit by issuing nationally backed bonds. Given that the raised amount is uncertain, participants can cover this uncertainty increasing the bond amount ratio. This adjustment is determined by the parameter $\kappa$, which can be any real number ranging from 1 to 2.5. More details are provided in Section 2.2.3.4.
year: 2025 water_utility: WU01 policies: bond_ratio: value: 2.0
2.3.2 Interventions
Interventions are physical modifications to the system, such as infrastructure upgrades or new installations. Interventions are specified annually and are always implemented. Note that if a utility contracts debt, a bond is issued automatically to cover the unbudgeted expenses.
2.3.2.1 Opening New Sources (Utility)
Participants can open new water sources to meet potential increases in demand. Available sources are predefined by location, and participants must specify a capacity within the allowable bounds.
year: 2025 water_utility: WU01 interventions: open_source: # Provide source identifier, capacity, and info about the connection - source_id: SG0158 source_capacity: 100 pump_option_id: PU003 n_pumps: 6 pipe_option_id: PI002 # Multiple sources can be added like this - source_id: SG0159 source_capacity: 50 pump_option_id: PU001 n_pumps: 4 pipe_option_id: PI003
2.3.2.2 Closing Sources (Utility)
Similarly, participants can close selected sources to improve the overall system efficiency.
year: 2025 water_utility: WU01 interventions: close_source: # Provide only the source identifier - source_id: SG0173 # Multiple sources can be removed like this. - source_id: SG0174
2.3.2.3 Installing Pipes (National or Utility)
Participants can decide to install new pipes or replace existing ones in the system. Each installation requires specifying the connection identifier and selecting a pipe option from a predefined list.
year: 2025 water_utility: WU01 interventions: install_pipe: # Provide connection identifier and pipe option - connection_id: CG0112 pipe_option_id: PI001 # Multiple pipes can be added like this. - connection_id: CG0113 pipe_option_id: PI008
2.3.2.4 Installing Pumps (Utility)
Participants can decide to install or replace pumps in pumping stations to calibrate the peak outflow of water sources. Each source is associated with one pumping station, which contains multiple identical pumps operating in parallel. Pump options must be selected from a predefined list. When installing pumps at an already-open pumping station, participants must specify whether to replace or add to the existing pumps. If the selected pump option differs from those already installed, all existing pumps will be automatically replaced, as only one pump type is allowed per station.
year: 2025 water_utility: WU01 interventions: install_pumps: # Provide source identifier, pump option and quantity for a new pumping station - source_id: SG0159 pump_option_id: PU003 n_pumps: 3 behaviour: replace # or "new"
2.3.2.5 Installing Solar (Utility)
Participants can decide to install behind-the-meter solar panels at water sources to reduce electricity costs and emissions for pumping stations. The solar panels offset electricity consumption and emissions but cannot be used as a profit-generating investment. Panels can be installed multiple times at different points in time. Note: Solar panels have a given lifespan (see Section 2.2.2); participants must decide whether to replace them upon expiration.
year: 2025 water_utility: WU01 interventions: install_solar: # Provide location and capacity - source_id: SG0158 capacity: 20 # Multiple sources can be added like this - source_id: SG0159 capacity: 20
@@ 2.4 System Requirements
2.4.1 Maintain Financial Viability
Participants must seek to maintain all water utilities in a financially viable state throughout the planning horizon. While utilities should not pursue profit, they must avoid insolvency.
Each year ($y$), the financial balance $F_{w}$ of every water utility $w$ is updated based on the following set of equations:
$F^{*}_{w}(y+1)$ $=F_{w}(y)+\text{NIB}(y)\cdot\alpha_{w}(y)+\text{REV}_{w}(y)$ (21)
$\quad-\text{CAPEX}_{w}(y)-\text{OPEX}_{w}(y)$
$\quad-\text{WLR}_{w}(y)-\text{WIC}_{w}(y)$
$\quad-\text{INT}_{w}(y)-\text{PRI}_{w}(y)$
[…]
$F_{w}(y+1)$ $=F_{w}^{*}(y+1)+\text{PRO}_{w}(y)$
The provisional fund balance $F_{w}^{*}(y+1)$ is calculated by accounting for all inflows and outflows:
  • $\text{NIB}(y)\cdot\alpha_{w}(y)$ is the national investment budget allocated for the water utility $w$ (see Section 2.3.1.1),
  • $\text{REV}_{w}(y)$ is the water utility’s revenue from the billable water demand (Equation 16),
  • $\text{CAPEX}_{w}(y)$ represents all the utility’s interventions capital costs (i.e., the sum of Equations 9, 10, 13, 14),
  • $\text{OPEX}_{w}(y)$ accounts for all the utility’s operational costs (i.e., the sum of Equations 8, 11),
  • $\text{WLR}_{w}(y)$ is the budget for NRW mitigation (see Section 2.3.1.2),
  • $\text{WIC}_{w}(y)$ is the cost for imported water from other water utilities (Equation 18),
  • $\text{INT}_{w}(y)$ the utility’s interest payments due (Equation 19), and
  • $\text{PRI}_{w}(y)$ is the principal amount due (Equation 19).
If the provisional balance is negative, the deficit defines the debt ($\text{debt}_{w}(y)$), which triggers bond issuance. The bond proceeds $\text{PRO}_{w}(y)$ are determined according to Section 2.2.3.4 and Equations 19, 20.
The actual fund balance $F_{w}(y+1)$ is obtained by adding the bond proceeds (if any) to the provisional balance, ensuring the fund remains solvent. Note that while surpluses are carried forward, deficits are always financed through bond issuance automatically ensuring that the fund balance is always positive.
Therefore, evaluation will focus only on the remaining debt at the end of the planning period. Final debt measures insolvency risk, not financial optimality. Competitors are free to pursue any financial strategy they deem appropriate, such as minimising costs, adjusting prices to enable expensive solutions, or adopting other innovative approaches. The many-objective framework ensures balanced evaluation across all dimensions. In other words, maintaining financial stability is an achievable target; the real challenge is deciding what to sacrifice.
2.4.2 Minimize GHG Emissions
Participants must keep the masterplan Greenhouse Gas (GHG) emissions to the lowest feasible level.
The GHG emissions for water utility $w$ in year $y$ are:
[…]
where $\mathrm{GHG}_{w}^{\text{emb}}(y)$ the embedded (construction) emissions, and $\mathrm{GHG}_{w}^{\text{op}}(y)$ the operational emissions from electricity use.
Only new pipes have embedded GHG emissions. The embedded emissions in year $y$ are:
[…]
where $\mathcal{C}_{w}$ is the set of connections of water utility ww444as described in Section 2.1.5, we distinguish between connections and pipes, $\mathbf{1}_{c\text{ activated in }y}$ is 1 if connection $c$ installs a new pipe in year $y$, 0 otherwise, $EF_{p_{c}}(y)$ is the unit emission factor of the connection’s selected pipe option $p_{c}$, and $L_{c}$ is the connection length.
The unit emission factor $EF_{p}$ for pipe option $p$ depends on the pipe option diameter $Diam_{p}$ and material $Mat_{p}$ and may change over time because of technological advancements, i.e., $EF_{p}(y)=EF(Diam_{p},Mat_{p},y)$.
The operational emissions are calculated based on the total electricity purchased from the grid, covering both water treatment (sources) and transport (pumping).
[…]
where for each timestep $t$ of a year yy555yy represent the year, while $\mathcal{Y}$ is the collection of timesteps, $E_{s}(t)$ and $EF_{s}(t)$ are the energy consumption and the emission factor of source $s$, while $E_{p}(t)$ and $EF_{p}(t)$ represent the same quantities for pump $p$.
Pumps energy consumption is retrieved via the EPANET simulations, while the water sources energy consumption is calculated according to Equation 7\$.
The emission factors of both entities (pumping stations $EF_{p}(t)$ and sources $EF_{s}(t)$) are dynamic and depend on the size and time of production of the behind-the-meter solar panels installation at that location (if no solar is installed, this variable reduces to the constant electricity grid emission factor in year $y$, i.e., $EF^{\text{el}}(y)$ ).
2.4.3 Maximize Service Reliability
Participants must ensure high service reliability by minimizing unmet water demand.
Service reliability for municipality $m$ in year $y$ is:
[…]
where $U_{m}(y)$ is the undelivered demand and $D^{\text{BIL}}_{m}(y)$ is the billable water demand.
Evaluation will focus on maintaining adequate service levels across all municipalities of each water utility throughout the entire planning horizon.
2.4.4 Promote Affordability and Equity
Participants must ensure water remains affordable, particularly for lower-income households, while maintaining equitable pricing across municipalities.
The affordability fairness metric represents the fraction of income that a household at the 20th percentile of the income distribution would spend on essential water consumption. Affordability fairness (lower is better) for water utility $w$ in year $y$ is:
[…]
where $P_{w}^{\text{fixed}}(y)$ and $P_{w}^{\text{variable}}(y)$ are the fixed and variable components of water price, $D^{\text{life}}$ is lifeline volume (minimum water required per person), and $ADI_{w}^{p20}(y)$ is the 20th percentile of disposable income across all households served by the water utility.
Evaluation will focus on minimizing affordability while maintaining reasonable equity across utilities.
@@ Acknowledgements
We thank all KWR colleagues for their invaluable suggestions and support throughout the project development. We would like to particularly thank Mirjam Blokker, Mark Morley, Karel van Laarhoven, Djordje Mitrovic, Ina Vertommen, Konstantinos Glynis, and Beatriz Gutierrez-Caloir.
Dennis, Lydia, Christos and Dragan would like to thank Yvonne Hassink-Mulder at Vitens for her insights on real water utility challenges; our collaboration profoundly shaped this battle from the early stages.
We would like to thank Luke Butler, whose tool EPANET.js was instrumental in developing this national network and enabled visualization of the information with unprecedented ease.
Dennis, Lydia and Christos would like to thank their academic supervisors Andrea Castelletti and Christos Makropoulos for their support and guidance.
D. Zanutto, C. Michalopoulos, L. Tsiami, J. Brandt, D. Eliades, S. Vrachimis, P. Samartzis, B. Hammer, P. Koundouri, M. Polycarpou, and D. Savić have received funding from the European Research Council (ERC) under the ERC Synergy Grant Water-Futures (Grant agreement No. 951424).
A. Artelt is supported by the Ministry of Culture and Science NRW (Germany) as part of the Lamarr Fellow Network.
This publication reflects the views of the authors only.
@@
All data versions are archived on Zenodo, with the latest version serving as the official standard (Zanutto et al. 2025). The files are also available in the releases section of the GitHub repository.
The model uses data from trusted open sources under Creative Commons licenses (e.g., CBS and KNMI) and draws on publicly available reports (e.g., Vewin). Where not explicitly cited, values are based on the authors’ institutional knowledge or represent modeling assumptions.
A complete description of all data sources will be released upon conclusion of the competition.
Notation
Entity Identifiers
Each entity is numbered sequentially starting from 1. Unique identifiers consist of two capital letters indicating the object type, followed by a number indicating the specific instance.
Jurisdictions:
  • Nation: NL0000
  • Region (Landsdeel): LD0000
  • Province: PV0000
  • Municipality (Gemeente): GM0000
Water Sources:
  • Groundwater: SG0000
  • Surface water: SS0000
  • Desalination: SD0000
Infrastructure:
  • Pumping stations: PS0000 (uniquely determines the source)
  • Pump option: PU000
Connections:
  • Between municipalities: CG0000
  • From source to municipality: CS0000
  • Between provinces: CP0000
  • Pipe option: PI000
Sets and Indices
[Our indexes here]
Parameters
[Our parameters here]
Decision Variables
[Our decision variables here]